OpenStudy (anonymous):

The slope of the line whose equation is 2x + 5y + 1 = 0 is

OpenStudy (kira_yamato):

5y + 2x + 1 = 0 ==> 5y = -2x - 1 ==> y = (-2x - 1)/5 Slope = -2/5

OpenStudy (across):

If you are lazy, then you can just perform implicit differentiation:\[2x+5y+1=0\]\[2+5y'=0\]\[y'=-2/5\]